# Metrics on completely positive maps via noncommutative geometry

### v1

@@ 3 Discussions on inducing metrics on unital completely positive maps
Note that if we have a sequence $\{L_{n}\}_{n\in\mathbb{N}}$ of seminorms $L_{n}\colon M_{n}(\mathbb{C})\otimes A\to[0,\infty]$, it is generally not true that $\{D_{L_{n}}\}_{n\in\mathbb{N}}$ will be a stable sequence of extended metrics. Even if $L_{n}=\|\cdot\|$, the operator norm on $M_{n}(\mathbb{C})\otimes A$, the sequence $\{D_{L_{n}}\}_{n\in\mathbb{N}}$ will not be stable. A way to fix this problem is as follows: Let
$D_{n}^{\rm stab}(F,G):=\sup\limits_{m\in\mathbb{N}}D_{L_{mn}}(\operatorname{id}_{mn}\otimes F,\operatorname{id}_{mn}\otimes G),$ (3.5)
where $L_{mn}$ is still the operator norm on $M_{mn}(\mathbb{C})\otimes A$. Then $\{D_{n}^{\rm stab}\}_{n\in\mathbb{N}}$ is stable. This is analogous to how the diamond norm, and the completely bounded norm are defined. Even though this sequence of metrics satisfies stability, it has the drawback that explicit calculations depend on calculating distances over all the matrix amplifications. Moreover, the construction only takes into consideration the geometric structure of $A$, not the target $C^{*}$-algebra $B$, therefore ignoring potentially important data present in the situation.
Another way of inducing a metric on $\operatorname{UCP}(A,B)$ discussed in [GLN05] and [BV25] involves embedding $\operatorname{UCP}(A,B)$ into the state space of a higher-dimensional $C^{*}$-algebra using the Choi-Jamiolkowski isomorphism. In Section 4 we consider a $C^{*}$-algebra $B$ with a faithful trace $\tau$, and we define the map $\omega_{\tau}\colon\operatorname{Hom}(A,B)\to\operatorname{Hom}(A\odot B^{\rm op},\mathbb{C})$ which gives an infinite-dimensional analogue of the map appearing in the Choi-Jamiolkowski isomorphism. We show that this gives an embedding $\omega_{\tau}^{\rm max}$ of $\operatorname{TC}_{\tau}(A,B)$ into $S(A\otimes_{\rm max}B^{\rm op})$. Furthermore, if $B$ is unital and $\tau$ is an amenable trace, then this gives an embedding $\omega_{\tau}^{\rm min}$ of $\operatorname{TC}_{\tau}(A,B)$ into $S(A\otimes_{\rm min}B^{\rm op})$, cf. Theorem 4.15. With this we can make the following definitions.

### v2

@@ 3. Discussions on inducing metrics on unital completely positive maps
We will study properties of such extended metrics in section 4.Note that if we have a sequence $\{L_{n}\}_{n\in\mathbb{N}}$ of seminorms $L_{n}\colon M_{n}(\mathbb{C})\otimes A\to[0,\infty]$, it is generally not true that $\{d_{L_{n}}\}_{n\in\mathbb{N}}$ will be a stable sequence of extended metrics. Even if $L_{n}=\|\cdot\|$, the $C^{*}$-norm on $M_{n}(\mathbb{C})\otimes A$, the sequence $\{d_{L_{n}}\}_{n\in\mathbb{N}}$ can not be expected to be stable (see [Wat05]). A way to obtain stability is to use the completely bounded norm $\|\cdot\|_{\rm cb}$ [Pau02], that is, to set
$d_{n}^{\rm stab}(F,G):=\|F-G\|_{\rm cb}=\sup\limits_{m\in\mathbb{N}}\|\operatorname{id}_{m}\otimes(F-G)\|,$
for all $F,G\in\operatorname{UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$. Then $\{d_{n}^{\rm stab}\}_{n\in\mathbb{N}}$ is stable. This is analogous to how the diamond norm on quantum channels is defined in the finite-dimensional case. Even though this sequence of metrics satisfies stability, the completely bounded metric has the drawback that explicit calculations depend on computing distances over all the matrix amplifications (except when $B$ is a matrix algebra, cf. [Smi83]). The same problem will occur with the stabilized version of $d_{L}$ introduced in section 4, which exists under natural assumptions. In fact, the computational complexity of $d_{L}$ itself is high as its definition requires taking two suprema.
As discussed in [GLN05] and [BV25], there is also a way of inducing a metric on the unital completely positive maps between matrix algebras by employing the Choi-Jamiołkowski isomorphism. By naively assuming that an embedding $\omega\colon\operatorname{UCP}(A,B)\to S(C)$ exists for some unital $C^{*}$-algebras $A,B$ and $C$, we could, given a seminorm $L$ on $C$, induce an extended metric on $\operatorname{UCP}(A,B)$ through
(3.5) $\Delta_{L}(F,G):=\operatorname{mk}_{L}(\omega(F),\omega(G)).$
In section 5, assuming that $B$ has a faithful trace, and choosing $C$ to be $A\otimes_{\rm max}B^{\rm op}$, or $A\otimes_{\rm min}B^{\rm op}$ under suitable assumptions, we develop the necessary technical tools to define an embedding $\omega$ to make the definition in (3.5) precise, see definition 6.1. In fact, this will allow us to characterize precisely the maps from $A$ to $B$ that will correspond to states on $A\otimes_{\rm max}B^{\rm op}$ under this embedding, cf. theorem 5.15.
4. Pullback-induced metrics
Let $A$ and $B$ be unital $C^{*}$-algebras, and $L\colon A\to[0,\infty]$ be a seminorm on $A$. We recall from definition 3.3 that $d_{L}\colon\operatorname{UCP}(A,B)\times\operatorname{UCP}(A,B)\to[0,\infty]$ is the extended metric given by
$d_{L}(F_{1},F_{2}):=\sup\limits_{\psi\in S(B)}\operatorname{mk}_{L}(F_{1}^{*}\psi,F_{2}^{*}\psi)=\sup\limits_{\psi\in S(B)}\operatorname{mk}_{L}(\psi\circ F_{1},\psi\circ F_{2})$
for all $F_{1},F_{2}\in\operatorname{UCP}(A,B)$. We will in this section establish sufficient conditions for a stabilized version of this metric to satisfy stability, see theorem 4.5, and chaining, see corollary 4.7.
A general problem with $d_{L}$ is that it is not clear under which conditions it is possible to define for each $n\in\mathbb{N}$ a seminorm $L_{n}$ on ${\rm UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$ such that $L_{1}=L$ and $\{d_{L_{n}}\}_{n\in\mathbb{N}}$ is stable. One way out of this problem is to use a stabilizing procedure. Before discussing this process, it is worth pointing out that $d_{L}$ is always stable in the following sense (employing product states):
Proposition 4.1.
Let $L:A\to[0,\infty]$ be a seminorm. For each $n\in\mathbb{N}$, let $1_{n}\otimes L$ be the seminorm on $M_{n}(\mathbb{C})\otimes A$ with ${\rm Dom}(1_{n}\otimes L)=M_{n}(\mathbb{C})\odot{\rm Dom}(L)$ given as in (2.3), that is,
$(1_{n}\otimes L)(x)=\sup\{L((\varphi\otimes\operatorname{id}_{A})(x))\colon\varphi\in S(M_{n}(\mathbb{C}))\}.$
Define $\delta_{1_{n}\otimes L}:{\rm UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)\times\operatorname{UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)\to[0,\infty]$ by
$\delta_{1_{n}\otimes L}(G_{1},G_{2}):=\sup\big\{\operatorname{mk}_{1_{n}\otimes L}\big((\varphi\otimes\psi)\circ G_{1},(\varphi\otimes\psi)\circ G_{2}\big):\varphi\in S(M_{n}(\mathbb{C})),\psi\in S(B)\big\}.$
Then each $\delta_{1_{n}\otimes L}$ is an extended metric on ${\rm UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$ satisfying that $\delta_{1_{n}\otimes L}\leq d_{1_{n}\otimes L}$ and $\delta_{1_{1}\otimes L}=d_{L}$. Moreover, $\{\delta_{1_{n}\otimes L}\}_{n\in\mathbb{N}}$ is stable.
Proof.
Let $n\in\mathbb{N}$. It is straightforward to verify that $\delta_{1_{n}\otimes L}$ is an extended pseudometric on ${\rm UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$ such that $\delta_{1_{n}\otimes L}\leq d_{1_{n}\otimes L}$. To show the separation property for $\delta_{1_{n}\otimes L}$, assume $G_{1}\neq G_{2}$ in ${\rm UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$. Pick then $x\in M_{n}(\mathbb{C})\otimes A$ such that $G_{1}(x)\neq G_{2}(x)$ in $M_{n}(\mathbb{C})\otimes B$. As the set of product states separates the points of the minimal tensor product of two unital $C^{*}$-algebras (see for instance [Val12] for a proof of this fact), there exist $\varphi\in S(M_{n}(\mathbb{C}))$ and $\psi\in S(B)$ such that
$(\varphi\otimes\psi)(G_{1}(x))\neq(\varphi\otimes\psi)(G_{2}(x)).$
By density of ${\rm Dom}(1_{n}\otimes L)$ in $M_{n}(\mathbb{C})\otimes A$, we can then find $y\in M_{n}(\mathbb{C})\otimes A$ such that $(1_{n}\otimes L)(y)<\infty$ and $|((\varphi\otimes\psi)\circ(G_{1}-G_{2}))(y)|>0$. Replacing $y$ with $\frac{1}{(1_{n}\otimes L)(y)}y$ when $(1_{n}\otimes L)(y)\neq 0$, we get that $(1_{n}\otimes L)(y)\leq 1$ and $|((\varphi\otimes\psi)\circ(G_{1}-G_{2}))(y)|>0$, hence that $\operatorname{mk}_{1_{n}\otimes L}((\varphi\otimes\psi)\circ G_{1},(\varphi\otimes\psi)\circ G_{2})>0$. This implies that $\delta_{1_{n}\otimes L}(G_{1},G_{2})>0$. Thus, $\delta_{1_{n}\otimes L}$ is an extended metric.
Next, it is not difficult to show that $\operatorname{mk}_{1_{n}\otimes L}\big(\varphi\otimes\psi_{1},\varphi\otimes\psi_{2}\big)=\operatorname{mk}_{L}(\psi_{1},\psi_{2})$ for every $\varphi\in S(M_{n}(\mathbb{C}))$ and $\psi_{1},\psi_{2}\in S(A)$. (We will actually prove a more general statement later, cf. lemma 6.3). Hence we get
$\delta_{1_{n}\otimes L}({\rm id}_{n}\otimes F_{1},{\rm id}_{n}$ $\otimes F_{2})$
$=\sup\{\operatorname{mk}_{1_{n}\otimes L}\big(\varphi\otimes(\psi\circ F_{1}),\varphi\otimes(\psi\circ F_{2})\big):\varphi\in S(M_{n}(\mathbb{C})),\psi\in S(B)\}$
$=\sup\{\operatorname{mk}_{L}\big(\psi\circ F_{1},\psi\circ F_{2}\big):\psi\in S(B)\}$
$=d_{L}(F_{1},F_{2})$
for all $F_{1},F_{2}\in{\rm UCP}(A,B)$. ∎
The process of producing a stabilized version of $d_{L}$ will rely on the following general result.
Proposition 4.2.
Assume that $d_{n}$ is an extended metric on ${\rm UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$ for each $n\in\mathbb{N}$, and that
(4.1) $d_{n}({\rm id}_{n}\otimes F_{1},{\rm id}_{n}\otimes F_{2})\leq d_{n+1}({\rm id}_{n+1}\otimes F_{1},{\rm id}_{n+1}\otimes F_{2})$
for all $n\in\mathbb{N}$ and $F_{1},F_{2}\in{\rm UCP}(A,B)$. For each $n\in\mathbb{N}$, set
(4.2) $d^{\,\rm stab}_{n}(G_{1},G_{2}):=\sup_{m\in\mathbb{N}}\big\{d_{mn}({\rm id}_{m}\otimes G_{1},{\rm id}_{m}\otimes G_{2})\big\}$
for all $G_{1},G_{2}\in{\rm UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$. Then we have that
$d_{1}^{\,\rm stab}(F_{1},F_{2}):=\lim_{m\to\infty}d_{m}({\rm id}_{m}\otimes F_{1},{\rm id}_{m}\otimes F_{2})$
for all $F_{1},F_{2}\in{\rm UCP}(A,B)$, and $\{d^{\,\rm stab}_{n}\}_{n\in\mathbb{N}}$ is a stable sequence of extended metrics.
Proof.
Note that in (4.2) we implicitly identify $M_{m}(\mathbb{C})\otimes M_{n}(\mathbb{C})$ with $M_{mn}(\mathbb{C})$, so that ${\rm id}_{m}\otimes G$ belongs to $\operatorname{UCP}(M_{mn}(\mathbb{C})\otimes A,M_{mn}(\mathbb{C})\otimes B)$ for every $G\in{\rm UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$.
The fact that $d^{\rm stab}_{n}$ is an extended metric on ${\rm UCP}(M_{n}(\mathbb{C})\otimes A,M_{n}(\mathbb{C})\otimes B)$ for every $n\in\mathbb{N}$ is easily verified. Let $n\in\mathbb{N}$ and $F_{1},F_{2}\in{\rm UCP}(A,B)$. Then the assumption (4.1) that the sequence $\{d_{n}({\rm id}_{n}\otimes F_{1},{\rm id}_{n}\otimes F_{2})\}_{n\in\mathbb{N}}$ is monotonically non-decreasing clearly implies that $d_{1}^{\,\rm stab}(F_{1},F_{2})=\lim_{k\to\infty}d_{k}({\rm id}_{k}\otimes F_{1},{\rm id}_{k}\otimes F_{2})$. Moreover, for $m,n\in\mathbb{N}$, (4.1) also implies that
$d_{mn}({\rm id}_{mn}\otimes F_{1},{\rm id}_{mn}\otimes F_{2})\leq d_{(m+1)n}({\rm id}_{(m+1)n}\otimes F_{1},{\rm id}_{(m+1)n}\otimes F_{2})$
for every $m\in\mathbb{N}$, i.e., the sequence $\{d_{mn}({\rm id}_{mn}\otimes F_{1},{\rm id}_{mn}\otimes F_{2})\}_{m\in\mathbb{N}}$ is monotonically non-decreasing. Thus we get
$d^{\,\rm stab}_{n}({\rm id}_{n}\otimes F_{1},{\rm id}_{n}\otimes F_{2})$ $=\sup_{m\in\mathbb{N}}\big\{d_{mn}({\rm id}_{m}\otimes{\rm id}_{n}\otimes F_{1},{\rm id}_{m}\otimes{\rm id}_{n}\otimes F_{2})\big\}$
$=\lim_{m\to\infty}d_{mn}({\rm id}_{mn}\otimes F_{1},{\rm id}_{mn}\otimes F_{2})$
$=\lim_{k\to\infty}d_{k}({\rm id}_{k}\otimes F_{1},{\rm id}_{k}\otimes F_{2})$
$=d_{1}^{\,\rm stab}(F_{1},F_{2}).\qed$
For $n\in\mathbb{N}$ let $\iota_{n}$ denote the canonical embedding of $M_{n}(\mathbb{C})$ into $M_{n+1}(\mathbb{C})$, determined by
$\iota_{n}\big(e_{i,j}^{(n)}\big)=e_{i,j}^{(n+1)}\quad\text{for all }1\leq i,j\leq n.$
The following definition will be helpful.